A ball is launched from ground level at 30 m/s at an angel of 35 degrees above the horizontal. How far does it go before it is atground level again?
A) 14 m
B) 21 m
C) 43 m
D) 86 m
E) 97 m
Answer #1
First determine the time:
y = Vit + 1/2at2
0 = 30m/s*sin(35)*t + 1/2*-9.81m/s2*t2
t = 3.51s
Then determine the distance:
x = V*t = 30m/s*cos(35)*3.51s = 86m